### v3

@@ 1 Introduction
In this work, we focus on the value of calibration for decision making. In the economics literature, the value of predictions is quantified by the payoff from decision making (Blackwell, 1951). Consider a decision maker who needs to choose an action $a\in A$ to maximize their decision payoff $U(a,\theta)$ , a (arbitrary) function of the action $a$ (e.g. take an umbrella or not) and the true binary state $\theta\in\{0,1\}$ (e.g. rainy or not). When the true state $\theta$ is unknown and only a prediction $p\in[0,1]$ is given, the decision maker can trust the prediction and take the action $a$ that maximizes the expected payoff $\operatorname{\mathbb{E}}_{\theta\sim p}[U(a,\theta)]$ , where “trust” means the decision maker assumes that the distribution of the state is indeed as predicted ($\Pr[\theta=1]=p$). Thus, the prediction manifests its value in the payoff received by the decision maker who best responds to the prediction as if the predictions were correct.
@@ 1.1 The Decision Loss from Miscalibration
Figure 1: In this example, ECE overestimates the decision loss from miscalibration. The plot visualizes the predictions in $[0,1]$. The best-response decision rule changes action at threshold $\nicefrac{{1}}{{2}}$ (red). The ground truth probability takes values from $0.2$ and $0.8$. The miscalibrated predictor predicts $0.4$ (blue) when the ground truth is $0.2$, and $0.6$ (blue) when the ground truth is $0.8$. Miscalibration induces no loss to the decision maker, since both the ground truth probability and the prediction recommends the same action.
@@
Similar to ECE, our CDL is a well-defined calibration error metric, with the basic property that a sequence of predictions is perfectly calibrated if and only if its CDL is zero, as implied by Foster and Vohra (1997). [^2] In the binary decision task in Figure 1, the CFDL of the miscalibrated predictor (Example 1.1) is 00. However, other decision tasks exist where the miscalibrated predictor has positive CFDL, leading to a non-zero CDL. However, we show that CDL is different from, and is not even a constant factor approximation of existing calibration error metrics, including ECE, the $K_{2}$ calibration error (i.e. the expected squared prediction bias), and the smooth calibration error (Błasiok et al., 2023a). In particular, we observe that ECE is often a significant overestimation of CDL. As a strong demonstration of this observation, we give an efficient algorithm for online prediction that minimizes CDL at a near-optimal rate of $O(T^{-1/2}\log T)$, which is faster than what is possible for ECE: it surpasses the $\Omega(T^{-0.472})$ lower bound for ECE by Qiao and Valiant (2021). We discuss this result with more details in Section 1.2 below.
@@ 1.2 Online Calibration
This remarkable first result of Foster and Vohra (1998) opened up the literature of online asymptotic calibration, with the focus of error metric mainly on ECE. On the upperbound side, subsequent work constructs polynomial-time algorithms for ECE minimization (e.g. Foster, 1999; Foster and Hart, 2021). On the lowerbound side, a recent breakthrough of Qiao and Valiant (2021) proves an $\Omega(T^{-0.472})$ lower bound for ECE, improving over the long-known natural $\Omega(T^{-1/2})$ bound. [^3] Subsequent to our work, Dagan et al. (2024) shows there exists a sequential prediction strategy that achieves strictly lower than $T^{-\nicefrac{{1}}{{3}}}$ expected ECE. They also give an improved lowerbound of ECE over $\Omega(T^{-0.472})$ in Qiao and Valiant (2021).
@@
Our result implies a substantial strengthening of an independent result by Roth and Shi (2024) on swap regret minimization. Swap regret minimization has been studied extensively in the literature of online learning (Hart and Mas-Colell, 2000; Blum and Mansour, 2007). Consider a fixed decision task repeated for $T$ rounds. The swap regret is the payoff increase (averaged over $T$ rounds) when the decision maker is allowed to apply a mapping $\sigma:A\to A$ and swap each action to another action in hindsight. The swap regret is stronger than the more commonly studied external regret, which only allows the decision maker to swap each action to a fixed action in hindsight. Foster and Vohra (1999) have observed that a predictor is perfectly calibrated if and only if it guarantees no swap regret for every downstream decision maker who trusts the prediction. Roth and Shi (2024) propose an algorithm that achieves $O(\frac{|A|}{\sqrt{T}})$ swap regret for every downstream decision maker, where $|A|$ is the number of actions.
@@
In our paper, our algorithm guarantees the same near-optimal $O(\frac{\log T}{\sqrt{T}})$ swap regret for every downstream decision maker simultaneously, removing the dependence on the number $|A|$ of actions in Roth and Shi (2024). This directly follows from the fact that swap regret is upperbounded by CFDL for a fixed a decision task. To see this, first recall that CFDL is the improvement in payoff when predictions are empirically calibrated to their conditional empirical frequencies. While empirically calibrated predictions may suggest different actions to the decision maker, this improvement can be equivalently written as the regret when, in hindsight, the decision maker is allowed to swap actions whenever they receive a different prediction. Thus, CFDL is stronger since the modification rule is finer-grained than that of the swap regret. If two predictions suggest the same best-response action to a decision maker, the modification rule in CFDL allows the decision maker to swap the actions differently, while the swap regret does not.
@@
The U-calibration error UCal, introduced by Kleinberg et al. (2023), is closely related to our CDL, despite being qualitatively weaker (i.e. UCal being zero is necessary but insufficient for calibration). Both errors are defined as the maximum payoff increase over decision tasks: CDL is the maximum payoff increase from calibrating the predictions; whereas U-calibration is the maximum payoff increase (a.k.a. external regret) from changing every prediction $p_{t}$ to the best fixed prediction (i.e., predicting the overall average $(\theta_{1}+\cdots+\theta_{T})/T$ every round). [^5] For a fixed decision task, the external regret may be negative, but the U-calibration error is non-negative because it is the maximum over all payoff-bounded decision tasks (including the degenerate task where the payoff is always zero). Therefore, the U-calibration error lowerbounds CDL. To see why UCal is not a well-defined calibration error, consider the miscalibrated predictor such that $(p_{t},\theta_{t})=(1/4,0)$ for $T/2$ rounds, and $(p_{t},\theta_{t})=(3/4,1)$ for the remaining $T/2$ rounds. The miscalibrated predictor always induce better or equal payoff compared to the best fixed prediction $1/2$, giving zero U-calibration error. See a detailed discussion in Example 4.5.
@@
While our result appears very similar to Kleinberg et al. (2023), we use substantially different techniques since our CDL is qualitatively stronger. Kleinberg et al. (2023) give a randomized algorithm that guarantees an optimal $O(T^{-1/2})$ expected U-calibration error. Their prediction algorithm ensures that each best-responding decision maker takes actions as if they are locally running the classic Hedge algorithm (see Arora et al., 2012), which is specific to the external regret. Thus, their algorithm does not extend to our CFDL.
@@ 1.3.1 Properties of CDL
We propose CDL as a calibration error metric and study its properties. In Section 3.3, we show CDL can be computed in time polynomial in the size of the prediction space $|Q|$ and the state space $|\Theta|$ by solving a linear program, which holds for general non-binary state space.
In Section 4, we discuss the separation between CDL and other calibration error metrics, including ECE, the $K_{2}$ calibration error, and the smooth calibration error. We consider the average calibration error calculated on $T$ samples of the prediction and the state. Our results show previous results on minimizing $K_{2}$ and smCal do not directly apply. We also show the separation between CDL and the U-calibration error UCal in Kleinberg et al. (2023), where $\textsc{UCal}=0$ is necessary but not sufficient for the predictions to be calibrated.
  • ECE is polynomially related to CDL.
@@
  • The smooth calibration error is not polynomially related to CDL, where we give examples.
@@ 1.3.2 Online CDL Minimization: Technical Overview
The key idea behind our $O(\frac{\log T}{\sqrt{T}})$ guarantee for CDL comes from the observation that the CDL can often be significantly smaller than ECE, despite their linear relationship in the worst-case. This allows us to bypass the $\Omega(\frac{T^{0.528}}{T})$ lower bound for ECE. Here we describe a typical example where the CDL is much smaller than ECE. Based on the intuition from this example, we establish a general lemma (Lemma 1.3) which plays a crucial role in our analysis.
Let us discretize the prediction space by restricting our predictor to only produce predictions in a finite set $Q:=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$, where $q_{i}=i/m$ for $i=1,\ldots,m$. We view each $q_{i}$ as a “bucket”, and thus our predictor makes predictions that fall into these buckets. For a sequence of $T$ predictions $\bm{p}=(p_{1},\ldots,p_{T})\in Q^{T}$ made by our predictor and the corresponding true states $\bm{\theta}=(\theta_{1},\ldots,\theta_{T})\in\{0,1\}^{T}$, let $n_{i}$ denote the number of predictions in bucket $i$:
@@
Thus, let us assume $|\widehat{q}_{i}-q_{i}|=T^{-1/4}$ for $i=1,\ldots,m$ for simplicity. Plugging this into (3), we get $\textsc{ECE}(\bm{p},\bm{\theta})=T^{-1/4}$. This is even worse than the $\textsc{ECE}=O(T^{-1/3})$ guarantee by the algorithm of Foster and Vohra (1998), mainly because here we discretize the prediction space into $m=\sqrt{T}$ buckets, but the optimal choice would be $m\approx T^{1/3}$.
However, the CDL in this example is much smaller: in fact we have $\textsc{CDL}\leq O(T^{-1/2})$. To prove this fact, let us first consider a specific decision task with action space $A=[0,1]$, where the payoff is given by
$U(a,\theta)=\mathbb{I}\left[\theta=\mathbb{I}\left[a>0.5\right]\right]=\begin{cases}\mathbb{I}\left[\theta=0\right],&\text{if $a\leq 0.5$};\\ \mathbb{I}\left[\theta=1\right],&\text{if $a>0.5$}.\end{cases}$
Here, actions in the interval $[0,0.5]$ are equivalent in terms of their payoffs, and similarly for actions in $(0.5,1]$. Therefore, the best response strategy is not unique. The best response strategy which we focus on here is the identity function: $a^{*}(p)=p$. The calibration fixed decision loss (CFDL) on decision task $U$ is defined as the improvement in payoff after the decision maker calibrates the predictor:
$\textsc{CFDL}_{U}(\bm{p},\bm{\theta})=\frac{1}{T}\sum_{t}\bigg{[}U(\sigma(p_{t}),\theta_{t})-U(p_{t},\theta_{t})\bigg{]},$ (4)
where $\sigma$ is the mapping that swaps each prediction to the empirical frequency on samples.
@@
If $q_{i}$ and $\widehat{q}_{i}$ belong to the same half of the interval $[0,1]$, i.e., $q_{i},\widehat{q}_{i}\in[0,0.5]$ or $q_{i},\widehat{q}_{i}\in(0.5,1]$, we clearly have
$\operatorname{\mathbb{E}}_{\theta\sim\widehat{q}_{i}}[U(\widehat{q}_{i},\theta)-U(q_{i},\theta)]=0.$ (6)
@@
$\operatorname{\mathbb{E}}_{\theta\sim\widehat{q}_{i}}[U(\widehat{q}_{i},\theta)-U(q_{i},\theta)]=\max(\widehat{q}_{i},1-\widehat{q}_{i})-\min(\widehat{q}_{i},1-\widehat{q}_{i})=2|\widehat{q}_{i}-0.5|\leq 2|q_{i}-\widehat{q}_{i}|.$ (7)
@@
$\textsc{CFDL}_{U}(\bm{a},\bm{\theta})\leq 2\sum_{i=1}^{m}n_{i}|q_{i}-\widehat{q}_{i}|\mathbb{I}\left[|q_{i}-0.5|\leq T^{-1/4}\right].$ (8)
Note that the number of $i$’s satisfying $|q_{i}-0.5|\leq T^{-1/4}$ is $O(T^{1/4})$. Thus, by our assumption of $n_{i}=\sqrt{T},|q_{i}-\widehat{q}_{i}|=T^{-1/4}$, we get $\textsc{CFDL}_{U}(\bm{a},\bm{\theta})=O(T^{-1/2})$.
While this bound on CFDL is for a specific decision task, we can extend it to every payoff-bounded decision task and get $\textsc{CDL}(\bm{p},\bm{\theta})=O(T^{-1/2})$. This follows from a result of Li ⓡ al. (2022) showing that any payoff-bounded decision task can be expressed as a convex combination of problems with V-shaped payoffs (see Section 3.2 for more details). Such decision tasks are all very similar to the one we consider here, and we can similarly obtain an $O(T^{-1/2})$ bound for the calibration loss for each of them. This reduction to V-shaped payoffs also played a crucial role in the $O(T^{-1/2})$ U-calibration guarantee of Kleinberg et al. (2023).
In our example above, we make the simplifying assumption that the number of predictions in each bucket is the same. However, the $O(T^{-1/2})$ bound on the CDL holds without this assumption while only losing a poly-logarithmic factor. That is, we have the following lemma:
Lemma 1.3 (Informal special case of Lemma 5.2 with $\alpha=O(1),\beta=0,\mathcal{D}(\bm{p},\bm{\theta})=0$).
@@
$\textsc{CDL}(\bm{p},\bm{\theta})\leq\widetilde{O}(T^{-1/2}).$
The intuition behind the lemma can be understood by analyzing the contribution of each bucket to CDL. In ECE, as expressed in Equation 3, the bias $n_{i}|q_{i}-\widehat{q}_{i}|$ in every bucket contributes to the sum, but in our example above, only a minority of the buckets make positive contribution to the CDL, as shown in (8). In general, for any specific decision task with V-shaped payoff, we show that the contribution to CDL from many buckets $i$ is significantly less than the bias $n_{i}|q_{i}-\widehat{q}_{i}|$. This allows us to prove the upper bound on CDL in Lemma 1.3, which would not hold if CDL were replaced by ECE.
Given Lemma 1.3, it remains to show that the guarantee $|\widehat{q}_{i}-q_{i}|\leq O(1/\sqrt{n}_{i})$ can indeed be (approximately) achieved in the online binary prediction setting. We use the result from Noarov et al. (2023) which, as stated, shows an efficient algorithm that gives us a bound only on the maximum of the expectation $\max_{i}({\mathbf{E}}\left[|\widehat{q}_{i}-q_{i}|\right]-O(1/\sqrt{n}_{i}))$, where the expectation is over the randomness in the algorithm. We refine their analysis and give a bound on the expectation of the maximum ${\mathbf{E}}\left[\max_{i}(|\widehat{q}_{i}-q_{i}|-O(1/\sqrt{n}_{i}))\right]$ (Lemma 5.5). As we show in Lemma 5.2 (a generalized version of Lemma 1.3), this bound is sufficient for us to obtain $\operatorname{\mathbb{E}}[\textsc{CDL}(\bm{p},\bm{\theta})]\leq O(\frac{\log T}{\sqrt{T}})$. We also note that a simpler minimax proof, similar to the one by Hart (2022), also allows us to show the existence of a randomized algorithm that approximately guarantees $|\widehat{q}_{i}-q_{i}|\leq O(1/\sqrt{n}_{i})$ and thus, by our Lemma 5.2 again, achieves $\operatorname{\mathbb{E}}[\textsc{CDL}(\bm{p},\bm{\theta})]\leq O(\frac{\log T}{\sqrt{T}})$ (see Appendix A). However, this proof does not come with an explicit construction or any efficiency guarantee.

### v4

@@ 1 Introduction
In this work, we focus on the value of calibration for decision making. In the economics literature, the value of predictions is quantified by the payoff from decision making (Blackwell, 1951). Consider a decision maker who needs to choose an action $a\in A$ to maximize their decision payoff $U(a,\theta)$, an arbitrary function of the action $a$ (e.g. take an umbrella or not) and the true binary state $\theta\in\{0,1\}$ (e.g. rainy or not). When the true state $\theta$ is unknown and only a prediction $p\in[0,1]$ is given, the decision maker can trust the prediction and take the action $a$ that maximizes the expected payoff $\operatorname{\mathbb{E}}_{\theta\sim p}[U(a,\theta)]$ , where “trust” means the decision maker assumes that the distribution of the state is indeed as predicted ($\Pr[\theta=1]=p$). Thus, the prediction manifests its value in the payoff received by the decision maker who best responds to the prediction as if the predictions were correct.
@@ 1.1 The Decision Loss from Miscalibration
Figure 1: In this example, ECE overestimates the decision loss from miscalibration for a specific decision task. The plot visualizes the predictions in $[0,1]$. The best-response decision rule changes action at threshold $\nicefrac{{1}}{{2}}$ (red). When the miscalibrated predictor predicts $0.4$ (blue), the actual empirical frequency is $0.2$; and when $0.6$ (blue) is predicted, the empirical frequency is $0.8$. Miscalibration induces no loss to the decision maker, since in both cases the prediction and the corresponding empirical frequency lie on the same side of the threshold, recommending the same action.
@@
Similar to ECE, our CDL is a well-defined calibration error metric, with the basic property that a sequence of predictions is perfectly calibrated if and only if its CDL is zero, as implied by Foster and Vohra (1997). [^2] In the binary decision task in Figure 1, the CFDL of the miscalibrated predictor (Example 1.1) is 00. However, other decision tasks exist where the miscalibrated predictor has positive CFDL, leading to a non-zero CDL. However, we show that CDL is different from, and is not even a constant factor approximation of existing calibration error metrics, including ECE, the $K_{2}$ calibration error (i.e. the expected squared prediction bias), the smooth calibration error, and the distance to calibration (Błasiok et al., 2023a). In particular, we observe that ECE is often a significant overestimation of CDL. As a strong demonstration of this observation, we give an efficient algorithm for online prediction that minimizes CDL at a near-optimal rate of $O(T^{-1/2}\log T)$, which is faster than what is possible for ECE: it surpasses the $\Omega(T^{-0.472})$ lower bound for ECE by Qiao and Valiant (2021). We discuss this result with more details in Section 1.2 below.
@@ 1.2 Online Calibration
This remarkable first result of Foster and Vohra (1998) opened up the literature of online asymptotic calibration, with the focus of error metric mainly on ECE. On the upperbound side, subsequent work constructs polynomial-time algorithms for ECE minimization (e.g. Foster, 1999; Foster and Hart, 2021). On the lowerbound side, a recent breakthrough of Qiao and Valiant (2021) proves an $\Omega(T^{-0.472})$ lower bound for ECE, improving over the long-known natural $\Omega(T^{-1/2})$ bound. [^3] Subsequent to our work, Dagan et al. (2024) show there exists a sequential prediction strategy that achieves $O(T^{-\nicefrac{{1}}{{3}}-\varepsilon})$ expected ECE for some $\varepsilon>0$. They also give an improved lowerbound for ECE over the $\Omega(T^{-0.472})$ in Qiao and Valiant (2021).
@@
Our result implies a substantial strengthening of an independent result by Roth and Shi (2024) on swap regret minimization. Swap regret minimization has been studied extensively in the literature of online learning (Hart and Mas-Colell, 2000; Blum and Mansour, 2007). Consider a fixed decision task repeated for $T$ rounds. The swap regret is the payoff increase (averaged over $T$ rounds) when the decision maker is allowed to apply a mapping $\sigma:A\to A$ and swap each action to another action in hindsight. The swap regret is stronger than the more commonly studied external regret, which only allows the decision maker to swap each action to a fixed action in hindsight. Foster and Vohra (1999) have observed that a predictor is perfectly calibrated if and only if it guarantees no swap regret for every downstream decision maker who trusts the prediction. Roth and Shi (2024) propose an algorithm that achieves $O\Big{(}|A|\sqrt{\frac{\log T}{T}}\Big{)}$ swap regret for every downstream decision maker, where $|A|$ is the number of actions.
@@
In our paper, our algorithm guarantees the same near-optimal $O(\frac{\log T}{\sqrt{T}})$ swap regret for every downstream decision maker simultaneously, removing the dependence on the number $|A|$ of actions in Roth and Shi (2024). This directly follows from the fact that swap regret is upperbounded by CFDL for a fixed a decision task. To see this, first recall that CFDL is the improvement in payoff when predictions are calibrated to the conditional empirical frequencies. While the calibrated predictions may suggest different actions to the decision maker, this improvement can be equivalently written as the regret when, in hindsight, the decision maker is allowed to swap actions whenever they receive a different prediction. Thus, CFDL is stronger since the modification rule is finer-grained than that of the swap regret. If two predictions suggest the same best-response action to a decision maker, the modification rule in CFDL allows the decision maker to swap the actions differently, while the swap regret does not.
@@
The U-calibration error UCal, introduced by Kleinberg et al. (2023), is closely related to our CDL, despite being qualitatively weaker (i.e. UCal being zero is necessary but insufficient for calibration). Both errors are defined as the maximum payoff increase over decision tasks: CDL is the maximum payoff increase from calibrating the predictions; whereas U-calibration is the maximum payoff increase (a.k.a. external regret) from changing every prediction $p_{t}$ to the best fixed prediction (i.e., predicting the overall average $(\theta_{1}+\cdots+\theta_{T})/T$ every round). [^5] For a fixed decision task, the external regret may be negative, but the U-calibration error is non-negative because it is the maximum over all payoff-bounded decision tasks (including the degenerate task where the payoff is always zero). Therefore, the U-calibration error lowerbounds CDL. To see why UCal is not a well-defined calibration error, consider the miscalibrated predictor such that $(p_{t},\theta_{t})=(1/4,0)$ for $T/2$ rounds, and $(p_{t},\theta_{t})=(3/4,1)$ for the remaining $T/2$ rounds. The miscalibrated predictor always induces better or equal payoff compared to the best fixed prediction $1/2$, giving zero U-calibration error. See a detailed discussion in Example 4.5.
@@
While our result appears very similar to the result of Kleinberg et al. (2023), we use substantially different techniques since our CDL is qualitatively stronger. Kleinberg et al. (2023) give a randomized algorithm that guarantees an optimal $O(T^{-1/2})$ expected U-calibration error. Their prediction algorithm ensures that each best-responding decision maker takes actions as if they are locally running the classic Hedge algorithm (see Arora et al., 2012), which is specific to the external regret. Thus, their guarantee does not directly extend to our CFDL.
@@ 1.3.1 Properties of CDL
We propose CDL as a calibration error metric and study its properties. In Section 3.3, we show, for general non-binary state space, CDL can be computed in time polynomial in the size of the prediction space $|Q|$ and the state space $|\Theta|$ by solving a linear program.
In Section 4, we discuss the separation between CDL and other calibration error metrics, including ECE, the $K_{2}$ calibration error, and the smooth calibration error. We consider the average calibration error calculated on $T$ samples of the prediction and the state. Our results show previous results on minimizing $K_{2}$ and smCal do not directly apply. We also show the separation between CDL and the U-calibration error UCal from Kleinberg et al. (2023), where $\textsc{UCal}=0$ is necessary but not sufficient for the predictions to be calibrated.
  • ECE is polynomially related [^6] If two error metrics $A$ and $B$ are polynomially related, we can find two polynomial functions of error $A$ as upperbound and lowerbound of error $B$, respectively. to CDL.
@@
  • The smooth calibration error and the distance to calibration are not polynomially related to CDL, where we give examples.
@@ 1.3.2 Online CDL Minimization: Technical Overview
In this section, we give an overview of our techniques for achieving near-optimal CDL. The key idea behind our $O(T^{-1/2}\log T)$ guarantee for CDL comes from the observation that the CDL can often be significantly smaller than ECE, despite their linear relationship in the worst-case. This allows us to bypass the $\Omega(T^{-0.472})$ lower bound for ECE. Here we provide a typical example where the CDL is much smaller than ECE. We will first describe the samples from a miscalibrated predictor, then show ECE is higher than CDL by calculation. Based on the intuition from this example, we establish a general lemma (Lemma 1.3) which plays a crucial role in our analysis.
Miscalibrated Predictor.
We now describe the observed samples from a miscalibrated predictor. The predictions are discretized to a finite set $Q:=\{q_{1},\ldots,q_{m}\}\subseteq[0,1]$, where $q_{i}=i/m$ for $i=1,\ldots,m$. We view each $q_{i}$ as a “bucket”, and thus our predictor makes predictions that fall into these buckets. We use $\mathbb{I}\left[\cdot\right]$ to denote the 0-1 indicator function: $\mathbb{I}\left[\text{statement}\right]=1$ if the statement is true, and $\mathbb{I}\left[\text{statement}\right]=0$ if the statement is false.
For a sequence of $T$ predictions $\bm{p}=(p_{1},\ldots,p_{T})\in Q^{T}$ made by our predictor and the corresponding true states $\bm{\theta}=(\theta_{1},\ldots,\theta_{T})\in\{0,1\}^{T}$, let $n_{i}$ denote the number of predictions in bucket $i$:
@@
Thus, in this example, we construct $|\widehat{q}_{i}-q_{i}|=T^{-1/4}$ for $i=1,\ldots,m$ for simplicity.
ECE.
ECE of the example above can be computed as
$\textsc{ECE}(\bm{p},\bm{\theta})=\frac{1}{T}\sum_{i=1}^{m}n_{i}|q_{i}-\widehat{q}_{i}|=T^{-1/4}.$ (3)
This is even worse than the $\textsc{ECE}=O(T^{-1/3})$ guarantee by the algorithm of Foster and Vohra (1998), mainly because here we discretize the prediction space into $m=\sqrt{T}$ buckets, but the optimal choice would be $m\approx T^{1/3}$.
CFDL and CDL.
However, the CDL in this example is much smaller: in fact we have $\textsc{CDL}\leq O(T^{-1/2})$. To prove this fact, let us first consider a specific decision task with action space $A=\{0,1\}$, where the payoff is given by
$U(a,\theta)=\mathbb{I}\left[\theta=a\right]\quad\text{for $a,\theta\in\{0,1\}$.}$
The best response strategy is to take a threshold at $0.5$:
$a^{*}(p)=\mathbb{I}\left[p>0.5\right]\quad\text{for $p\in[0,1]$}.$
The calibration fixed decision loss (CFDL) on decision task $U$ is defined as the improvement in payoff after the decision maker calibrates the predictor:
$\textsc{CFDL}_{U}(\bm{p},\bm{\theta})=\frac{1}{T}\sum_{t=1}^{T}\bigg{[}U(a^{*}(\sigma(p_{t})),\theta_{t})-U(a^{*}(p_{t}),\theta_{t})\bigg{]},$ (4)
where $\sigma$ is the mapping that moves each prediction to the actual empirical frequency.
@@
If $q_{i}$ and $\widehat{q}_{i}$ belong to the same half of the interval $[0,1]$, i.e., $q_{i},\widehat{q}_{i}\in[0,0.5]$ or $q_{i},\widehat{q}_{i}\in(0.5,1]$, we have $a^{*}(\widehat{q}_{i})=a^{*}(q_{i})$ and thus
$\operatorname{\mathbb{E}}_{\theta\sim\widehat{q}_{i}}[U(a^{*}(\widehat{q}_{i}),\theta)-U(a^{*}(q_{i}),\theta)]=0.$ (6)
@@
$\operatorname{\mathbb{E}}_{\theta\sim\widehat{q}_{i}}[U(a^{*}(\widehat{q}_{i}),\theta)-U(a^{*}(q_{i}),\theta)]$ $=\operatorname{\mathbb{E}}_{\theta\sim\widehat{q}_{i}}[U(a^{*}(\widehat{q}_{i}),\theta)]-\operatorname{\mathbb{E}}_{\theta\sim\widehat{q}_{i}}[U(a^{*}(q_{i}),\theta)]$
$=\max(\widehat{q}_{i},1-\widehat{q}_{i})-\min(\widehat{q}_{i},1-\widehat{q}_{i})$
$=2|\widehat{q}_{i}-0.5|$
$\leq 2|q_{i}-\widehat{q}_{i}|.$ (7)
@@
$\textsc{CFDL}_{U}(\bm{a},\bm{\theta})\leq\frac{2}{T}\sum_{i=1}^{m}n_{i}|q_{i}-\widehat{q}_{i}|\mathbb{I}\left[|q_{i}-0.5|\leq T^{-1/4}\right].$ (8)
Note that the number of $i$’s satisfying $|q_{i}-0.5|\leq T^{-1/4}$ is $O(T^{1/4})$. Thus, by our assumption of $n_{i}=\sqrt{T}$ and $|q_{i}-\widehat{q}_{i}|=T^{-1/4}$, we get $\textsc{CFDL}_{U}(\bm{a},\bm{\theta})=O(T^{-1/2})$.
While this bound above on CFDL is for a specific decision task, we can extend it to every payoff-bounded decision task and get $\textsc{CDL}(\bm{p},\bm{\theta})=O(T^{-1/2})$. This follows from a result of Li ⓡ al. (2022) showing that any payoff-bounded decision task can be expressed as a convex combination of tasks with V-shaped payoffs (see Section 3.2 for more details). Such decision tasks are all very similar to the one we consider here, and we can similarly obtain an $O(T^{-1/2})$ bound for the calibration loss for each of them. This reduction to V-shaped payoffs also played a crucial role in the $O(T^{-1/2})$ U-calibration guarantee of Kleinberg et al. (2023).
Intuition Generalized.
In our example above, we make the simplifying assumption that the number of predictions in each bucket is the same. However, the $O(T^{-1/2})$ bound on the CDL holds without this assumption while only losing a logarithmic factor. That is, we have the following lemma:
Lemma 1.3 (Informal special case of Lemma 5.2).
@@
$\textsc{CDL}(\bm{p},\bm{\theta})\leq O(T^{-1/2}\log T).$
The intuition behind the lemma can be understood by analyzing the contribution of each bucket to CDL. In ECE, as expressed in Equation 3, the bias $n_{i}|q_{i}-\widehat{q}_{i}|$ in every bucket contributes to the average, but in our example above, only a minority of the buckets make positive contribution to the CDL, as shown in (8). In general, for any specific decision task with V-shaped payoff, we show that the contribution to CDL from many buckets $i$ is significantly less than the bias $n_{i}|q_{i}-\widehat{q}_{i}|$. This allows us to prove the upper bound on CDL in Lemma 1.3, which would not hold if CDL were replaced by ECE.
Given Lemma 1.3, it remains to show that the guarantee $|\widehat{q}_{i}-q_{i}|\leq O(1/\sqrt{n}_{i})$ can indeed be (approximately) achieved in the online binary prediction setting. We use the result from Noarov et al. (2023) which, as stated, shows an efficient algorithm that gives us a bound only on the maximum of the expectation $\max_{i}({\mathbf{E}}\left[|\widehat{q}_{i}-q_{i}|\right]-O(1/\sqrt{n}_{i}))$, where the expectation is over the randomness in the algorithm. We refine their analysis and give a bound on the expectation of the maximum ${\mathbf{E}}\left[\max_{i}(|\widehat{q}_{i}-q_{i}|-O(1/\sqrt{n}_{i}))\right]$ (Lemma 5.5). As we show in Lemma 5.2 (a generalized version of Lemma 1.3), this bound is sufficient for us to obtain $\operatorname{\mathbb{E}}[\textsc{CDL}(\bm{p},\bm{\theta})]\leq O(\frac{\log T}{\sqrt{T}})$. We also note that a simpler minimax proof, similar to the one by Hart (2022), also allows us to show the existence of a randomized algorithm that approximately guarantees $|\widehat{q}_{i}-q_{i}|\leq O(1/\sqrt{n}_{i})$ and thus, by our Lemma 5.2 again, achieves $\operatorname{\mathbb{E}}[\textsc{CDL}(\bm{p},\bm{\theta})]\leq O(\frac{\log T}{\sqrt{T}})$ (see Appendix A). However, this proof does not come with an explicit construction or any computational efficiency guarantee.